OpenStudy (anonymous):

Find the indicated product. Leave the answer in polar form. 6(cos38degrees+isin38degrees)*3(cos62degrees+isin62degrees). a. 9(cos24° + isin24°) b.18(cos100° + isin100°) c. 9(cos100° + isin100°) d. 18(cos24° - isin24°)

jimthompson5910 (jim_thompson5910):

hint: (r1*cis(theta1))*(r2*cis(theta2)) = (r1*r2)*cis(theta1+theta2)

OpenStudy (anonymous):

let me give you an example: 5(cos(25)+isin(25))*(3(cos(10)+isin(10))=15(cos(35)+isin(35)).

OpenStudy (anonymous):

Thanks! Is it b.?

jimthompson5910 (jim_thompson5910):

it is